An oxide of iron consists of 70 %by mass of iron and the rest is oxygen.find it's simplest formula
1
Expert's answer
2020-03-31T12:32:49-0400

mass % of oxygen ="100 \%-70\%=30 \%"

Let the mass of iron oxide to be 100 g, then m(Fe) = 70 g, m(O)=30 g

"n(Fe):n(O)=\frac{70g}{55.85g/mol}:\frac{30g}{16.00g/mol}= 1.253:1.875"


"\frac{1.253}{1.253}:\frac{1.875}{1.253} = 1:1.5 = 2:3"


The simplest formula is "Fe_2O_3"